# Automated Static Vulnerability Detection via a Holistic Neuro-symbolic Approach

### v1

@@ 5. Implementation
Instrumentation of Symbolic Validator.Instrumentation of Symbolic ValidatorInstrumentation of Symbolic Validator. We realized the runtime program state tracking by instrumenting the query execution runtime. In particular, we enhanced the query parsing process to locate syntax errors. For the other two, we hooked the internal language interpreter. Specifically, these query languages’ interpreters typically contain a main interpretation loop that switches over the instruction types and invokes specific handlers. We thus added additional hookers for selected handlers to collect variables and their runtime values.
@@
Query Execution Server.Query Execution ServerQuery Execution Server. The iterative query generation would invoke the symbolic validator and static analysis runtime multiple times. The invocations often operate on the same codebase—the project containing the input vulnerability example. To improve the overall efficiency, we eliminate the repetitive project loading process by starting a local server to host the codebase for interactive, continuous query execution. This could significantly improve the efficiency of the symbolic validator.
@@
Function References in Joern.Function References in JoernFunction References in Joern. During our implementation, we realized that Joern could not correctly connect a function’s reference to its definition. When generating queries, the LLMs (and human experts) all would first assume such function inferences are properly handled. This introduces a lot of failures in the generated queries. Therefore, we carefully modified the variable references in Joern (Yamaguchi et al., 2014) and complemented the DSL of Joern to support explicit function dereference.
6. Evaluation
In this section, we extensively evaluate MoCQ to answer the following questions.
@@ 6. Evaluation
    New Vulnerability Detection.New Vulnerability DetectionNew Vulnerability Detection. How effective is MoCQ in discovering new vulnerabilities in real-world applications?
@@
    Query Effectiveness.Query EffectivenessQuery Effectiveness. How effective are the MoCQ-generated queries for detecting vulnerabilities? How do they compare to related approaches?
@@
    Ablation Study.Ablation StudyAblation Study. How does each component of MoCQ contribute to its performance?
[…]
    DSL Subsetting.DSL SubsettingDSL Subsetting. What are the characteristics of the DSL subset used by MoCQ?
6.1. Experimental Setup
[…]
Dataset.DatasetDataset. We consider seven types of vulnerabilities in PHP and JavaScript from OWASP Top Ten (OWASP Foundation, 2021) for our evaluation. For each vulnerability type, we selected 20 popular projects that contain known vulnerabilities. We examined relevant public sources, including the CVE database and GitHub issues, and incorporated findings from our previous research, to collect all known vulnerabilities as the ground truth on a best-effort basis. In total, we included $265$ vulnerabilities for our evaluation. Vulnerabilities within a single project may share similarities, potentially leading to data leakage if some are used for query generation and others for performance evaluation. To mitigate this, we randomly selected 15 projects (192 vulnerabilities) for query generation—the generation dataset, and the remaining 5 projects (73 vulnerabilities) for performance evaluation—the testing dataset. We consistently adhere to this dataset split throughout the entire evaluation. In addition to the two datasets with ground truth, we collect 19 projects in their latest versions to evaluate new vulnerability discovery—referred to as the latest project dataset.
[…]
Configuration of MoCQ.Configuration of MoCQConfiguration of MoCQ. MoCQ is a complex system influenced by various factors, particularly the number of projects and vulnerabilities used for query generation, as well as the choice of underlying model and static analysis engine. Unless otherwise noted, our default setting involves generating queries from vulnerability examples drawn from ten randomly selected projects in the generation dataset, using the Claude 3.7 Sonnet model and static analysis engine Joern. These ten projects cover 113 input vulnerability examples. A more comprehensive ablation study evaluating other settings is presented in § 6.4. We ran the experiments three times with an attempt threshold of $5,000$ iterations and used the mean values for the results.
6.2. Zero-day Detection
We first evaluate MoCQ on the latest project dataset. MoCQ-generated queries effectively discover seven zero-day vulnerabilities, shown in Table 1, including four JavaScript prototype pollution, one JavaScript command injection, one PHP SQL injection, and one PHP cross-site scripting. We responsibly reported the vulnerabilities and to date one has been acknowledged.
As a comparison, we collected queries from the official Joern tool (joe, 2025) as well as from recent publications (Li and Meng, 2021; Fröberg, 2023). When there are multiple queries for a vulnerability type, we use the one with the best detection results. On the same dataset, expert-crafted queries identified three new vulnerabilities, all of which were also detected by MoCQ-generated queries. The other four vulnerabilities were exclusively detected using the new patterns discovered by MoCQ. Our in-depth study revealed four unique vulnerability patterns that contributed to MoCQ’s superior performance; these were entirely missed by existing Joern queries, leading to their failure to detect the vulnerabilities. A more comprehensive list of new patterns uncovered by MoCQ is discussed in § 6.3.2 and Table 4.
Table 1. New vulnerabilities and the contributing new patterns.
Project Vul. Type Experts’ MoCQ Contributing New Pattern Status
forkcms (N/A, 2025d) PHP SQLi - Rep.
SyliusCmsPlugin (N/A, 2025a) PHP XSS $\times$ Ctx-sensitive san. via htmlspecialchars Rep.
parse-static-imports (N/A, 2025e) JS Cmd. $\times$ Vulnerable operation with repl Ack.
es2015-proxy (N/A, 2025c) JS Proto. $\times$ Data flow via Object.assign Rep.
web-worker (N/A, 2025g) JS Proto. - Rep.
koa-send (N/A, 2025h) JS Proto. - Rep.
content-type (N/A, 2025b) JS Proto. $\times$ Data flow via Reflect.set Rep.
[…]
Case Study.Case StudyCase Study. Figure 4 shows the prototype pollution vulnerability in es2015-proxy (N/A, 2025c). The mergeOptions function iterates over the properties of source object and assigns the property values to target object. When source is crafted to include the special property "__proto__", the mergeOptions function inadvertently assigns its value to overwrite target["__proto__"], which refers to the object’s prototype. Specifically, the proof-of-concept (PoC) payload we created modifies the implementation of the toString method of target["__proto__"]. This modification impacts not only target but also all objects that inherit from the foundational object prototype Object.prototype, such as the newObj on line 14. Detecting this vulnerability requires capturing the Object.assign in the pattern for property modification, which was missed by expert-crafted queries.
1function mergeOptions(target, source) {2    for (let key in source) {3      Object.assign(target[key], source[key]);4    }5    return target;6}78// Proof-of-Concept (PoC)9const obj = {}10const payload = ’{"__proto__": {"toString": "Hacked"}}’;11mergeOptions(obj, JSON.parse(payload));1213// All objects have the polluted property14const newObj = {};15console.log(newObj.toString); // "Hacked"
Figure 4. A JavaScript prototype pollution vulnerability.
[…]
MoCQ Experts’ Detection Delta MoCQ+Experts’ Pure LLM
[…]
PHP SQLi 19 16 14 0.84 0.53 17 21 0.89 0.45 15 1 2 18 23 0.95 0.44 9 45 0.47 0.17
PHP XSS 13 8 10 0.62 0.44 7 14 0.54 0.33 7 1 0 8 15 0.62 0.35 9 27 0.69 0.25
PHP Type. 6 4 22 0.67 0.15 3 18 0.50 0.14 3 1 0 4 24 0.67 0.14 2 23 0.33 0.08
PHP Deser. 8 5 4 0.62 0.56 4 6 0.50 0.40 2 3 2 7 7 0.88 0.50 5 57 0.63 0.08
JS Proto. 11 10 8 0.91 0.56 7 12 0.64 0.37 6 4 1 11 13 1.00 0.46 4 39 0.36 0.09
JS Cmd. 5 5 17 1.00 0.23 5 15 1.00 0.25 5 0 0 5 20 1.00 0.20 3 25 0.60 0.11
JS XSS 11 8 15 0.73 0.35 9 11 0.82 0.45 6 2 3 11 12 1.00 0.48 6 34 0.55 0.15
Total / *Avg. 73 56 90 *0.77 *0.40 52 97 *0.70 *0.34 44 12 8 64 114 *0.87 *0.37 38 250 *0.52 *0.13
6.3. Query Effectiveness
We apply the queries generated under the default setting to the testing dataset and compare them to related approaches.
6.3.1. Performance of MoCQ
@@
The testing dataset contains 72 true vulnerabilities and per-type breakdown is shown as TPtotal in Table 2. Overall, MoCQ-generated queries could detect a broad range of vulnerabilities across PHP and JavaScript ecosystems. Out of 72 known vulnerabilities, MoCQ-generated queries successfully detected 56 cases with 90 false positives. We also computed the recall ($Rec.=\frac{TP}{TP+FN}=\frac{TP}{TP_{total}}$) and precision ($Prec.=\frac{TP}{TP+FP}$). MoCQ achieved a strong overall recall of 0.77 and a precision of 0.40. The precision, especially the false positive rate, is reasonable in practice, especially given the challenge of statically analyzing dynamic PHP and JavaScript codebases. As we will show in § 6.3.2, even expert-crafted queries have similar precision.
@@
We analyzed the 16 vulnerabilities missed by MoCQ-generated queries. These false negatives stem from multiple root causes. There are 11 cases caused by incomplete patterns, including missing taint source specifications—particularly involving customized third-party sources—missing vulnerable operations, etc., which could have been detected if a more comprehensive pattern had been generated. Besides, two cases involved incomplete data flows where the targets of dynamic function calls were missing from the project’s program representations. This occurred because the static construction of representations failed to properly handle dynamic language features. Three cases were caused by flaws in type-related issues. For instance, inaccurate type information led to one false negative in PHP type juggling and two in JavaScript prototype pollution. The latter two categories could not be fixed by refining the patterns.
6.3.2. Comparison to Expert-crafted Queries
We further evaluated the collected expert-crafted queries on the same testing dataset.
Table 3. Query generation efficiency.
PHP SQLi PHP XSS PHP Type. PHP Deser. JS Proto. JS Cmd. JS XSS
MoCQ
# Examples 23 28 11 13 15 10 13
Time 15.3h 24.6h 5.2h 18.7h 21.4h 7.9h 10.5h
Experts’
# Commits 5 5 10 17 28 1 1
Time Span 2d 3d 7d 5w 7w - -
[…]
Engineering Efforts.Engineering EffortsEngineering Efforts. MoCQ is efficient in generating queries and demonstrates a significant advantage in required engineering efforts. We report the number of input vulnerability examples and the total query generation time per vulnerability type in Table 3. In general, MoCQ required only a few hours to generate queries for each vulnerability type, significantly shorter than previous best practices that relied heavily on expert engineering. Generally, types with more input examples took longer to process. On average, it took 0.9 hours to construct a query for a single example across different types. Complex vulnerabilities like PHP deserialization and JavaScript prototype pollution were the most time-consuming cases, requiring 1.4 and 1.5 hours per example, respectively.
We analyzed the Git commit history in the open-source repositories of the expert-crafted queries. We find that experts often use multiple commits to develop and revise the queries, and the commit history often spans days to a few weeks. We acknowledge that Git commit history does not precisely reflect the actual, continuous engineering effort involved. For instance, the time span between the first and last commit may overestimate or underestimate the true effort, as substantial work may have occurred before the first commit or between commits without being recorded. Nevertheless, we argue that it serves as a reasonable proxy for estimating the engineering effort required. As a comparison, MoCQ, as a fully-automated solution, can run continuously and shorten the amount of time to construct queries.
[…]
In the remaining four types, however, MoCQ-generated and expert-crafted queries demonstrated complementary strengths—each was able to detect some vulnerabilities missed by the other. Motivated by this, we explored how their combination would perform. We combined both MoCQ-generated and experts’ queries, and then used the combined set for Joern in vulnerability detection (MoCQ+Experts’). This combined version achieved promising results, with a high recall of 0.87 and a precision of 0.37 on average. Notably, it successfully detected all known vulnerabilities in three vulnerability types, achieving a perfect recall of 1. This shows that although MoCQ-generated queries alone do not surpass expert-crafted queries in all cases, their combination is an effective strategy.
Following that, we tried to apply MoCQ to refine the expert-crafted query for JavaScript prototype pollution as an example. MoCQ ran for around one hour and automatically revised the query, which in return detected 3 previously missed vulnerabilities. This highlights that MoCQ not only automatically generates detection queries but also can improve an existing expert-crafted query.
[…]
New Patterns.New PatternsNew Patterns. We highlight the new patterns MoCQ discovered. We define an atomic operation—such as a specific API call, expression type, or data flow—that is not recognized by expert-crafted queries as a new pattern. We then conservatively locate these new operations through pair-wise comparison with expert-crafted queries and found 12 new vulnerability patterns, as shown in Table 4. We constructed a minimal PoC vulnerability example to validate each new pattern. Compared to expert-crafted queries, MoCQ uncovered new operations that propagate data flows, or new vulnerability operations that can lead to security consequences. By filling in these missing components, static analysis could more comprehensively search vulnerabilities in target application code.
MoCQ not only found nine new patterns specific to a vulnerability type but also three general patterns applicable to the vulnerabilities in the entire programming language. For example, four mechanisms for PHP type juggling were missed by experts, e.g., PHP built-in functions array_search and array_flip where implicit type casting could occur. MoCQ also detected a new general way in JavaScript to propagate data flows via Reflect.set—which could benefit the detection of different types of vulnerabilities. We submitted the new patterns to the maintainers of expert-crafted queries. One of them has been acknowledged.
Table 4. New vulnerability patterns uncovered by MoCQ. *MoCQ uncovered 34 distinct operations in the phar category for PHP deserialization and we conservatively counted one.
Vul. Type Pattern Functionality
General PHP htmlspecialchars Ctx-sensitive sanitization
PHP Type. in_array Implicit type casting
PHP Type. array_search Loose comparison
PHP Type. array_flip Implicit type casting
PHP Type. case in switch Type coercion in case matching
PHP Deser. phar* Object deserialization
PHP Deser. copy Object deserialization
General JS Object.assign Data propagation
General JS Reflect.set Data manipulation
JS Proto. Object.assign Data propagation
JS Proto. Object.defineProperty New property manipulation
JS Proto. obj.__defineGetter__ New property manipulation
6.3.3. Comparison to LLM-based Approaches
We compare MoCQ to LLM-based approaches including pure LLM-based and IRIS (Li et al., 2025).
[…]
@@ 6.3.3. Comparison to LLM-based Approaches
Pure LLM-based Detection.Pure LLM-based DetectionPure LLM-based Detection. This approach directly provides the source code to an LLM and asks it to assess if there is any vulnerability. Unlike MoCQ, there is no LLM-based tool that can perform whole project analysis for large codebases, to the best of our knowledge. We thus take the common practice by separately checking individual code snippets (source code file) (Ding et al., 2025; Mohajer et al., 2024) with Claude 3.7 Sonnet. To handle non-determinism across multiple LLM runs, we assign the final label based on majority agreement—if at least two out of three runs agree, we adopt that label for the case.
@@
The evaluation results in the last part of Table 2 show lower precision and recall. This is primarily due to the lack of cross-file information, such as global variables and inter-procedural function calls defined outside the current file, leading to many false positives and false negatives. For instance, detecting JavaScript prototype pollution requires tracking data flow from external user input to object property manipulation. It is thus often infeasible when analyzing a single file in isolation. We also observed that when the entire vulnerability is contained within a single file, the LLM tends to produce fairly reliable results.
[…]
@@
IRIS.IRISIRIS. IRIS (Li et al., 2025) is a related work that combines LLM-identified taint specifications with existing queries. We consider its contribution orthogonal or complementary to that of MoCQ. IRIS is designed for Java programs and is not directly applicable for an end-to-end comparison. We thus incorporated its main technique and evaluated how could help resolve our false negatives. The results showed that IRIS could successfully identify missing data sources for three false negative cases. Nevertheless, IRIS does not have a holistic, fully-automated way of query generation.
6.4. Ablation Study

### v3

@@ 5.2. RQ1: New Vulnerabilities and Patterns
PHP type. 6 4 22 0.67 0.15 3 18 0.50 0.14 4 24 0.67 0.14 2 23 0.33 0.08 7d 5.2h 14m 20m
@@
JS proto. 11 10 8 0.91 0.56 7 12 0.64 0.37 11 13 1.00 0.46 4 39 0.36 0.09 7w 21.4h 19m 26m
@@
JS XSS 11 8 15 0.73 0.35 9 11 0.82 0.45 11 12 1.00 0.48 6 34 0.55 0.15 - 10.5h 20m 27m
@@
C/C++ heap ov. 4 3 6 0.75 0.33 3 2 0.75 0.60 3 5 0.75 0.38 1 18 0.25 0.05 4w 9.4h 11m 14m
@@
Java SQLi 3 3 5 1.00 0.38 2 2 0.67 0.50 3 3 1.00 0.50 1 15 0.33 0.06 3w 8.5h 12m 16m
@@
Overall 94 73 119 0.79 0.39 65 111 0.67 0.40 81 137 0.85 0.39 45 341 0.44 0.11 - 160.1h 183m 260m
5.3. RQ2: Detection Effectiveness
We evaluate MoCQ on D2 and compare to related approaches.
5.3.1. Detection Results of MoCQ
@@
D2 contains true vulnerabilities and per-type breakdown is shown as $TP_{total}$ in Table 3. Overall, MoCQ-generated queries could detect a broad range of vulnerabilities across PHP, JavaScript, C/C++, and Java ecosystems. Out of 94 known vulnerabilities, MoCQ-generated queries successfully detected 73 cases with 119 false positives. We also computed the recall ($\frac{TP}{TP_{total}}$) and precision ($\frac{TP}{TP+FP}$). MoCQ achieved a strong overall recall of 0.79 and a precision of 0.39. The precision, especially the false positive rate, is reasonable in practice, particularly for statically analyzing dynamic languages like PHP (Luo et al., 2022) and JavaScript (Shcherbakov et al., 2023) and unmanaged languages like C/C++ (Sui and Xue, 2016). As we will show in § 5.3.2, expert-developed queries have similar precision.
@@
We analyzed the 21 vulnerabilities missed by MoCQ-generated queries. These false negatives stem from multiple root causes. There are 14 cases caused by incomplete patterns, including missing taint source specifications—particularly involving customized third-party sources—missing vulnerable operations, etc., which could have been detected if a more comprehensive pattern had been generated. Besides, four cases involved incomplete data flows where the targets of dynamic function calls were missing from the project’s program representations. This occurred because the static construction of representations failed to properly handle dynamic language features and indirect calls. Three cases were caused by flaws in type-related issues. For instance, inaccurate type information led to one false negative in PHP type juggling and two in JavaScript prototype pollution. The latter two categories could not be fixed by refining the patterns.
5.3.2. Comparison to Expert-developed Queries
We evaluated the collected expert-developed queries on the same dataset to thoroughly assess the effectiveness. Overall, MoCQ-generated queries achieved comparable performance to expert-developed queries, despite the latter requiring substantial security expertise and engineering effort. As shown in Table 3, expert-developed queries detected slightly fewer true positives (65 vs. 73) and produced slightly fewer false positives (111 vs. 119), resulting in an average recall of 0.67 and a precision of 0.40. A detailed comparison reveals that MoCQ and expert-developed queries detected 54 vulnerabilities in common, indicating a substantial overlap in their detection capabilities. In seven vulnerability types (PHP XSS, PHP type juggling, JavaScript command injection, C/C++ UAF, C/C++ heap overflow, Java SQLi, Java API misuse), MoCQ-generated queries detected all vulnerabilities found by the expert-developed queries.
However, our in-depth manual analysis of the queries reveals the different cognitive strengths of LLMs and human experts. In particular, LLMs excel at capturing subtle low-level linguistic nuances that humans might overlook, while human experts are better at high-level conceptual reasoning and understanding broader vulnerability contexts. For example, expert-developed queries all overlooked PHP’s case-insensitivity (e.g., echo is the same as eCHo), which was detected by MoCQ’s strength in syntactic nuances. Conversely, expert queries showed superior design in identifying complex multi-step SQL injection (Dahse and Holz, 2014) that require understanding broader application context.
This observation motivated us to investigate whether combining both sets of queries could lead to improved detection. We applied Joern using the union of MoCQ-generated and expert-developed queries, resulting in a combined approach that achieved a high recall of 0.85 and a precision of 0.39 on average. These results highlight that while MoCQ alone may not always outperform expert-developed queries, combining them yields a robust and highly effective detection strategy. Following that, we conducted a preliminary experiment to apply MoCQ to refine the expert-developed query for JavaScript prototype pollution as an example. MoCQ ran for around one hour and automatically revised the query, which in return detected 3 previously missed vulnerabilities. This highlights that MoCQ can not only automatically generate detection queries but also improve an existing expert-developed query.
5.3.3. Comparison to Other Approaches
We compare MoCQ to other related approaches on D2. In particular, we include a diverse set of methods, including a pure LLM-based approach, a prior neuro-symbolic method, and state-of-the-art static methods.
@@ 5.3.3. Comparison to Other Approaches
Pure LLM-based Detection. This approach directly provides the source code to an LLM and asks it to assess if there is any vulnerability. Unlike MoCQ, there is no LLM-based tool that can perform whole project analysis for large codebases, to the best of our knowledge. We thus take the common practice by separately checking individual code snippets (source code file) (Ding et al., 2025; Mohajer et al., 2024) with Claude 3.7 Sonnet. To handle non-determinism across multiple LLM runs, we assign the final label based on majority agreement—if at least two out of three runs agree, we adopt that label for the case.
@@
The evaluation results in the last part of Table 3 show lower precision and recall. This is primarily due to the lack of cross-file information, such as global variables and inter-procedural function calls defined outside the current file. This leads to many false positives and false negatives. For instance, detecting JavaScript prototype pollution requires tracking data flow from external user input to object property manipulation. It is thus often infeasible when analyzing a single file in isolation. Providing more context to the LLM may potentially improve the results, however, LLMs inherently have the scalability issue for large-scale codebases.
@@
IRIS. IRIS (Li et al., 2025) is a related work that combines LLM-identified taint specifications with existing queries. We consider its contribution orthogonal or complementary to that of MoCQ. IRIS is not directly applicable to the evaluated languages or vulnerability types for an end-to-end comparison. We thus incorporated its main technique and evaluated how it could help resolve our false negatives. The results showed that IRIS could successfully identify missing data sources for three false negative cases. Nevertheless, IRIS does not have a holistic, automated way of synthesizing end-to-end patterns.
Other State-of-the-art Static Methods. Our comparison to expert-developed queries demonstrates MoCQ’s advantage of automation. We are also interested in how MoCQ performs compared to other general, state-of-the-art, non-query-based static analysis tools. However, these tools often specialize in single vulnerability types, unlike MoCQ’s broad multi-language, multi-vulnerability support. It is non-trivial and time-consuming to set up each specialized tool for all covered vulnerability types. We thus select three representative, state-of-the-art analyzers (LChecker (Li and Meng, 2021) for PHP type juggling, Silent-Spring (Li et al., 2022) for JavaScript prototype pollution, and Cred (Yan et al., 2018) for C/C++ UAF) and evaluated them on D2. These tools employ different design and implementation choices (e.g., LLVM-based, CFG-based). We thus believe they effectively represent the diversity and sophistication of modern static analysis paradigms. Unlike MoCQ, all these approaches require a lot of engineering of vulnerability patterns.
LChecker reported six true positives and 37 false positives, while MoCQ reported four true positives and 22 false positives. LChecker benefits from a more precise type inference algorithm than the built-in type information available in Joern, which helps improve recall. However, in dynamic languages like PHP, many variable types often cannot be resolved statically. As a result, LChecker conservatively assigns multiple possible types, which can lead to over-approximation and more false positives. A more robust type inference would help Joern and MoCQ improve both precision and recall. For prototype pollution detection, Silent-Spring reported ten true positives and 12 false positives, achieving the same recall as MoCQ (91%) but a lower precision of 45%. For UAF, Cred reported four true positives with fewer false positives (four). This is because Cred leverages LLVM-based pointer analysis, which enables more accurate resolution of memory access patterns in C/C++ programs than Joern’s graph-based analysis. Nevertheless, MoCQ represents a significant step in automating pattern-based analysis.
5.4. RQ3: Performance and Cost
We measure the performance and cost of MoCQ, and present the results in the last four columns of Table 3.
Query Generation Time. As shown in Table 3, MoCQ required only a few hours to generate queries for each vulnerability type, significantly shorter than previous best practices that relied heavily on expert engineering. Generally, types with more input examples took longer to process. On average, it took 0.9 hours to construct a query on a single vulnerability example across different types.
We analyzed the Git commit history in the open-source repositories of the expert-developed queries. We find that experts often use multiple commits to develop and revise the queries, and the commit history often spans days to a few weeks. We acknowledge that Git commit history does not precisely reflect the actual, continuous engineering effort involved. For instance, the time span between the first and last commit may overestimate or underestimate the true effort, as substantial work may have occurred before the first commit or between commits without being recorded. Nevertheless, it serves as a reasonable proxy for comparing development timelines. As a comparison, MoCQ, as an automated solution, can run continuously and shorten the amount of time to construct queries.
Analysis Time. As shown in the last two columns of Table 3, MoCQ’s automated queries took 260 minutes in total to complete analysis on the tested projects, moderately longer than the 183 minutes required by expert-developed queries. This difference arises because most MoCQ patterns are broader and thus take more time to execute. Analysis time depends on application complexity and query design, but the overall duration of MoCQ’s remains practical for real-world use.
API Cost. Query generation per vulnerability type under the default setting (Claude Sonnet 3.7) typically takes less than 150 iterations with costs \$36 USD on average for LLM inference. Once generated, these queries can scan unlimited codebases without additional API cost.
5.5. RQ4: Ablation Study